A mixture of argon and carbon dioxide gases at a total pressure of 866 mm Hg contains argon at a partial pressure of 507 mm Hg. If the gas mixture contains 7.53 grams of argon, how many grams of carbon dioxide are present?
Answer #1

Solution. Given that :- Tote Pressure of gases(P)= 366 mmHg Partial pressure of Argon Par) =507 MMHG mass of Argon (m) = 7.530.5855(0-1885 molethcox) = 01885 mole 0.1104 +0.5855 nco, = 0.1885 mole 0.5855 No, = 6.1885-001104) mole 0.5855860, 0:0781 monco2 = 0.1334 mole molar mass of 202=449/mole so mass of CO2 = 0.1334 molex 44 g ) 5,879 mole mass of CO2 = 5.87 g Ans.